OpenStudy (grant330sims):

Find the coordinates of the vertex for the parabola defined by the given quadratic function: f(x)=x^2+2x-8

OpenStudy (anonymous):

Given function is a parabola of the standard form: -A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex. A is a multiplier which affects the steepness or slope of the parabola. The minus sign of the lead coefficient means that the parabola opens downwards, that is, it has a maximum. .. Finding the coordinates of the vertex: f(x) = -x^2 -4x + 8 completing the square y=-(x^2+4x+4)+8+4 y=-(x+2)^2+12 Ans: Coordinates of the vertex: (-2,12)

OpenStudy (grant330sims):

well, i know thats not right, I know the x is -1, because of -b/2a. But after that, dont you just put the x in to find y?